write an equation whose solution representsfollowing statement.

a. The value that when increased by 6 is 17.
b.The value that when decreased by 37 is 74.
c. The value that is 2 smaller that 10.

Answer by Theo(13342)

a. x + 6 = 17

b. x - 37 = 75

c. x + 2 = 10

x + 6 = 17 becomes x = 11 which is 6 smaller than 17.
when you increase 11 by 6 you get 17.

x - 37 = 75 becomes x = 112.
when you decrease 112 by 37, you get 75.

x + 2 = 10 becomes x = 8

8 is smaller than 10 by 2.